OpenStudy (anonymous):

[4_07]Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. y = -x + 6 y = x - 6 y = -x - 2 y = x + 2

OpenStudy (anonymous):

are you even typeing?

OpenStudy (anonymous):

I was, I accidentally did the wrong thing, sorry!

OpenStudy (anonymous):

it's ok just the same thing was with the last one

OpenStudy (anonymous):

if it perpendicular, m1m2 = -1 m1=-1 -1/-1=1 so, m2=1 if you then substitute the coordinates (-4,-2) into the two equations with 1 as the gradient you get: y = x - 6 -2 = -4 - 6 -2=-10 this is incorrect, so it's not this equation. try it with the second: y = x + 2 -2 = -4 + 2 -2=-2 this is correct, so y = x + 2 is the correct equation :)

OpenStudy (anonymous):

thank you